The data required for a new investment are calculated as follows:

 Cost of the investment: 10.000.000 $

 Annual growth rate expected from cash flows is 10%

 Cost of capital (discount rate): 25%

 Economic life of the investment: 5 years

 Tax rate: 40%

 Normal depreciation method is applied.

 If the firm is expected of investment every year for 5 years, the depreciation and pre-tax profit (AVOK) will be 7,000,000 TL, how much will the net current value of investment will be?


1
Expert's answer
2021-04-15T07:33:23-0400

Profit before tax and depreciation =7000000

Annual depreciation =10000000/5

Annual depreciation =2000000

Profit after depreciation =70000000-2000000=5000000

Profit after tax =5000000(1-0.4)

Profit after tax =3000000

PAT=3000000

Present value of growing ANNUITY =P((1+r)5-(1+g)5)/(r-g)

Present value of ANNUITY =3000000((1.25)5-(1.1)5)/(0.25-0.1)

Value of firm =28824956.25TL